How do I use transformations of y=x^4 or y=x^5 to graph these functions.
1) f(x)=(x+3)^5
2) f(x)= x^5+3
3) f(x)= (x-1)^4-2



Answer by stanbon(75887)

How do I use transformations of y=x^4 or y=x^5 to graph these functions.
1) f(x)=(x+3)^5
Shift y = x^5 three units to the left.
---------------------------------------------
2) f(x)= x^5+3
Raise each point of y = x^5 three units.
---------------------------------------------
3) f(x)= (x-1)^4-2
Shift y = x^4 one unit to the right; then lower each point two units.
-------------------------
Cheers,
Stan H.